OpenStudy (foolaroundmath):

If an amount of $200 is compounded at a rate of 5% annually, what is the interest paid after 2.5 years. I know the formula is \(P *(1 + r/n)^{nT} - P\). Here P is 200 r = 0.05 n = 1 and T = 2.5 ?? Is it permitted to have fractional T? Can you explain or give some sources where it is explained why T can/ can't be fractional.

OpenStudy (misty1212):

T is time in years, if you are compounding annually then T should be a whole number

OpenStudy (foolaroundmath):

Hmm, then what about in this particular question? Will T be 2.5? Btw, the original question is from a previous year CFA question paper.

OpenStudy (foolaroundmath):

@misty1212

OpenStudy (foolaroundmath):

@satellite73 @dan815 @shrutipande9

OpenStudy (amistre64):

how many months is half a year?

OpenStudy (amistre64):

ugh, how many years is .5 years is what i meant to ask lol

OpenStudy (amistre64):

B (1+r) (1+r) (1+r/2) yr1 yr2 yr.5 youve only aquired half the interest for the partial year

OpenStudy (foolaroundmath):

@amistre64 , so the calculation would be \[P (1 + r)^{2}(1+\frac{r}{2})^{0.5}\] Am I right?

OpenStudy (amistre64):

not the .5 exponent that only occurs once, not 1/2 a time

OpenStudy (foolaroundmath):

I see what you're saying. So it would just be \[P(1+r)^{2}(1+\frac{r}{2}) \] So basically we treat the half year as being compounded half-yearly

OpenStudy (freethinker):

\(n \times T = 1 \times 2.5\)

OpenStudy (amistre64):

now what im reading is telling me that\[B(1+r)^{k}~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~(1)\] "in the mathematical theory of interest, if we say that an account earns compound interest at a rate i, we are implicitly stating that we use formula (1) for partial periods as well" https://www.math.purdue.edu/~rcp/MA170/InterestTheory.pdf in practice, institutions tend to just do simple interest for the partial year it says as well so the thing is, if this is for a course, what does your course work give?

OpenStudy (amistre64):

B(1+r)^2 is the begnining balance, and if we close out our account, we receive the extra simple interest for the partial period. B(1+r)^2 * (1 + r/2)

OpenStudy (foolaroundmath):

The book doesn't explicitly state anything for partial periods. Thanks for the link though.

OpenStudy (amistre64):

page 4 starts talking of your issue here :)

OpenStudy (foolaroundmath):

Yep found it :) Thanks for the help.

OpenStudy (amistre64):

youre welcome